It is given that 


z+2i=iz+λ, (w/z)=2+2i, Imw=8


where z and w are complex numbers, and λ is a real constant. Which of the following is the value of λ? 

  1. λ=4
  2. λ=8
  3. λ=3
  4. λ=-3
  5. λ=-4
1
Expert's answer
2021-10-21T11:53:00-0400

Let "z=a+bi, w=c+8i." Then


"a+bi+2i=i(a+bi)+\lambda, \lambda\in \R"

"b+2=a"

"a=-b+\lambda"

"b=a-2"

"\lambda=a+a-2"

"(w/z)=\dfrac{c+8i}{a+bi}=2+2i"

"c+8i=2a+2bi+2ai-2b"

Then


"c=2a-2b"

"8=2b+2a"

"a+b=4"

Substitute


"a+a-2=4=>a=3"

"\lambda=3+3-2=4"

5. "\lambda=4"